OpenStudy (anonymous):

I am trying to figure out 200P50 I know that I use my calculator which is 200 nPr 50 but it says overflow is there a simpler way of doing this? (permutations)

OpenStudy (ddcamp):

When finding nPr, you can use the equation: \[\frac{ n! }{ (n-r)! }\] to find the permutations.

OpenStudy (anonymous):

i keep getting overflow error...

OpenStudy (sweetburger):

just wondering what level math is this?

OpenStudy (anonymous):

I am in college problem solving math.... I hear it is stuff you are supposed to know from high school

OpenStudy (sweetburger):

oh god im in 10th grade precal and i have never seen it before

OpenStudy (ddcamp):

\[\frac{ 200! }{ (200-50)! }=\frac{ 200! }{ 150! } =\frac{ 200\times199\times...\times151\times150\times149\times...\times3\times2\times1 }{ 150\times149\times...\times3\times2\times1 }\\ = 200\times199\times...\times152\times151\]

OpenStudy (anonymous):

do you think that leaving it as 200!/150! will be a good enough answer?